Investment. Thomas inherited $10,000 and invested it in two treasury bonds which pay 4% and 5% simple interest. The annual interest is $435. Find the amount invested at 5%?
Answer by stanbon(75887)

Investment. Thomas inherited $10,000 and invested it in two treasury bonds which pay 4% and 5% simple interest. The annual interest is $435. Find the amount invested at 5%?
------------------
Equations:
Quantity: x + y = 10000
Interest:4x + 5y = 43500
--------
Multiply Quantity by 4:
4x + 4y = 40000
4x + 5y = 43500
---
Subtract and solve for "y":
y = $3500 (amt. invested at 5%)
=======================
Cheers,
Stan H.
===================